a. A restaurant manager wishes to calculate the minimum sales point (MSP) when considering whether they should remain open one extra hour per day. Which will NOT be required for them to compute to make the proper MSP calculation?

A. Fixed cost %
B. Variable cost %
C. Food cost %
D. Minimum payroll costs for the time period

b. Which business cost will increase as a restaurant's volume increases?

A. Mortgage costs
B. Insurance costs
C. Napkin costs
D. Lighting costs

c. A variable cost is one that...

A. increases as volume increases and decreases as volume decreases.
B. decreases as volume increases and decreases as volume decreases.
C. is unaffected by changes in volume.
D. increases as volume increases and increases as volume decreases.

d. A manager is considering staying open one extra hour per day and wants to calculate the minimum sales point (MSP). The manager's variable costs are 40%, food costs are 30%, and fixed costs are 60%. The minimum labor cost for staying open one extra hour is $125.00. Rounded up to the nearest dollar, what is the manager's MSP?

A. $457
B. $417
C. $407
D. $437
Answer #1
a ANSWER :
A. Fixed Cost %
Fixed cost is fixed whether the restaurant is open for one more hour or not
Hence , it is not relevant to calculate MSP
b Napkin cost will increase with increase in volume
Mortgage Cost, Insurance Cost and Lighting Costs are fixed costs
ANSWER: C.Napkin Costs
c Variable Cost is one that :
A.increases as volume increases and decreases as volume decreases.
Variable cost like cost of materials varies with the volume of output
d Assume MSP=X
Managers Variable =0.4X
Food Cost =0.3X
Labor cost for one hour extra=$125
Total Cost for one our extra opening :
o.4X+0.3X+125=0.7X+125
For Break even,
X=Total cost for one hour extra opening
X=0.7X+125
(1-0.7)X=125
X=125/0.3= $417
ANSWER:B. $417